1. Amy made semiannual deposits of $3,100 at the beginning of every every six months into a fund earning 6.8% compounded semiannually for nine years. No further deposits were made.

   a. How much will be in the account 15 years after the first deposit? _____________    

   b. How much in total was deposited? ______________

c. How much interest will Amy earn? _____________

 

 

2. What is the principal invested at 4.75% compounded semiannually from which monthly withdrawals of $240 can be made at the beginning of each month for 15 years? _________________

 

 Need it in 7 hours

 

 


 

 

 

 




1
Expert's answer
2021-03-22T08:25:18-0400

1.a) In the account after 9 years:

"FV_9=(R(1+i)^n-1)\/i"

"FV_9=(3100\cdot1.068^{9\cdot2}-1)/0.068=\$148967.82"

In the account after 9 years:

"FV_{15}=FV_9(1+i)^n"

"FV_{15}=148967.82\cdot1.068^{6\cdot2}=\$328055.63"


b) It was deposited:

"P=3100\cdot9\cdot2=\$55800"


c) Amy will earn:

"FV_{15}-P=328055.63-55800=\$272255.63"



2.

"P_0=d(1-(1+i)^{-n})\/i"

"P_0=240(1-(1+0.0475/6)^{-15\cdot12})/0.0475/6=\$22983.56"